The probability that a student uses the Academic Resource Center on a regular basis is 0.42 . In a group of 20 students, what is the probability that exactly 7 of them use the Academic Resource Center on a regular basis? Write only a number as your answer. Round to 4 decimal places (for example 0.2416). Do not write as a percentage.
Answer #1

Solution

Given that ,

p = 0.42

1 - p = 1 - 0.42 = 0.58

n = 20

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X = 7 ) = ( 20C 7) * (0.42)7 * (0.58)13

= 0.150218

Probability = 0.1502